A Syracuse University freshman’s diet consists of pizza and hot dogs. The price of one pizza is $5 and one hot-dog is $2. At the beginning of the semester, the pizza restaurant on campus decides to run the following promotion: If a student brings in the sticker on the pizza box back to the restaurant, he gets a refund of $1 for every sticker beyond the first 10 stickers that he returns. Assume that a freshman has $100 to spend on pizza and hot dogs during the semester. Draw a typical freshman’s budget line. Make sure to carefully label your graph and all the relevant points.
Answer #1

Budget line:

M = Px*x +Py *y

X =pizza

Y = Hot dog

100 = 5*x+2*y

For x>10 income = 100+(x-10)

New BL

5*x+2*y =100+(x-10)

4*x + 2*y = 90

The budget line is:

  

     

5*x+2*y =100 for x<10

  4*x + 2*y = 90 for x>10